# Complexity at Scale: A Quantitative Analysis of an Alibaba Microservice Deployment

### v2

@@ 7. Future Work
In this paper, we have undertaken an extensive quantitative exploration into a large-scale MSA at Alibaba. Still, because of the substantial amount of data produced by large-scale MSAs, this investigation is not exhaustive and raises open questions that warrant future investigation. For instance, we found that the degree of MSs varies in time (Section 5.1), however, we did not investigate whether other local dependency features change, nor if the global topology changes, in time. Our analysis of call graph variability (Section 4.7) does not account for call graph structure, potentially underestimating true variability. Our findings indicated that the MSs with the most total calls, number of instances, and dependencies (Section 4.5) were generally not the same, however, there are other aspects of the dataset that may help elucidate links between these groups. To encourage further exploration from the community on this dataset, we have released all code and processed versions of the dataset (Winchester, 2025).
@@
Appendix A Ethics
This work does not raise any ethical issues.

### v3

@@ 8. Future Work
In this paper, we have undertaken an extensive quantitative exploration into a large-scale MSA at Alibaba. Still, because of the substantial amount of data produced by large-scale MSAs, this investigation is not exhaustive and raises open questions that warrant future investigation. For instance, we found that the degree of MSs varies in time (Section 6.1), however, we did not investigate whether other local dependency features change, nor if the global topology changes, in time. Our analysis of call graph variability (Section 5.6) does not account for call graph structure, potentially underestimating true variability. Our findings indicated that the MSs with the most total calls, number of instances, and dependencies (Section 5.4) were generally not the same, however, there are other aspects of the dataset that may help elucidate links between these groups. To encourage further exploration of this dataset by the research community and to ensure replicability, we will release all code and processed data used in our analysis upon acceptance of this paper.